OpenStudy (anonymous):

Subtract five and one-fourth minus three and five-eighths. two and one-eighth two and five-eighths one and one-eighth one and five-eighths

OpenStudy (dumbcow):

\[\frac{1}{4} - \frac{5}{8} = \frac{2}{8} - \frac{5}{8} = -\frac{3}{8}\] so 5-3 = 2 put it together \[2-\frac{3}{8} = 1 \frac{8}{8} -\frac{3}{8} = 1 \frac{5}{8}\]

OpenStudy (anonymous):

Ty so much @dumbcow